A car is traveling around a horizontal circular track with radius r 270 m at a constant speed v 19 m/s as show angle θA-25° above the x axis, and the angle θ8 66 below the x axis 1) What is the magnitude of the cars acceleration? m/s2 Submit You currently have 2 submissions for this question. Only 10 submission are allowed You can make 8 more submissions for this question 2) What is the x component of the cars acceleratiorn when it is at point A /s2 Submit You currently have 2 submissions for this question. Only 10 submission are allowed. You can make 8 more submissions for this question. 3) What is the y component of the cars acceleration when it is at point A m/s2 Submit You currently have O submissions for this question. Only 10 submission are allowed. You can make 10 more submissions for this question. 16 MacBook Air


media%2F9e7%2F9e7e369d-a8b0-49ef-97b2-40
Answer #1

r = 270 m ; v =19m/s ; \theta_{A}=25^{\circ}; \theta_{B}=66^{\circ}

(1)

a =\frac{v^{2}}{r}=\frac{(19)^{2}}{270}=1.33m/s^{2}

Keep in mind that the acceleration vector points toward the origin, so add 180º to the given angles.

(2)

a_{xA}=a*cos(180+25)^{\circ}=1.33* cos205^{\circ}=-1.025m/s^{2}

(3)

a_{yA}=a*sin(180+25)^{\circ}=1.33* sin205^{\circ}=-0.562m/s^{2}

(4)

a_{xB}=a*cos(180-66)^{\circ}=1.33* cos114^{\circ}=-0.54m/s^{2}

(5)

a_{yB}=a*sin(180-66)^{\circ}=1.33* sin114^{\circ}=1.215m/s^{2}